# Reputation in the Shadow of Exit

### v4

@@ 5.2 Selecting Equilibria
When is it that the threat of cruelty is enough to deter revolution? Since the game is monotone and satisfies the conditions of Corollary 1, asymptotic $c^{*}$-incentive compatibility is both necessary and sufficient for the regime to deter revolution (as much as possible) in the unique Markovian equilibrium. Rewriting the incentive constraints to the simpler setting of the global game described above yields that asymptotic $c^{*}$-incentive compatibility is equivalent to the inequality
$\delta[V^{*}(\delta)-\underline{V}_{0}(\delta)]>\kappa(c^{*})$
holding for all $\delta$ large and the optimal deterrent choice $c^{*}$. In the global games setting, asymptotic $c^{*}$ incentive compatibility is related to the degree of dispersion of higher order beliefs around the risk-dominant threshold at the average, i.e. whether $\mu>(<)\frac{1}{1+c^{*}}$. Here, recall value $\frac{1}{1+c^{*}}$ is the risk-dominant signal threshold selected by the global game as signals become arbitrarily precise, i.e. $\lim\limits_{\varsigma\to 0}x^{*}(c^{*})=\frac{1}{1+c^{*}}$ by Frankel et al. (2003).
What is the right commitment posture $c^{*}$ to consider when comparing to $\mu$? There are two concerns. First, since $c^{*}(\delta,\varsigma)$ varies with the discount rate and noise precision, the risk-dominant selection threshold varies with the discount rate as well. Second, the underlying probability distributions $(F_{\varepsilon},P_{\theta})$ will also affect the optimal choice for the regime and affect their commitment payoff. To resolve this problem, consider the set
$c^{*}(1,0)=\limsup_{\delta\to 1}\left(\limsup_{\varsigma\to 0}c^{*}(\delta,\varsigma)\right)$
where $\limsup A_{n}$ are the subsequential limits of all selections from $A_{n}$. That $c^{*}(1,0)$ is nonempty follows from two compactness arguments, first by noting $c^{*}(\delta,0)$ is well-defined and nonempty for each $\delta$ and hence so much $c^{*}(1,0)$. Each element of $c^{*}(1,0)$ is approximately optimal for all $\delta\geq\bar{\delta}$ close enough to $1$ and $\varsigma<\bar{\varsigma}(\delta)$, depending potentially on choice of $\delta$, close enough to 0. Throughout, I will fix some $c^{*}\in c(1,0)$ and appeal to continuity of the problem in parameters to make the argument that the same statement holds for $c^{*}(\delta,\varsigma)$ when the first parameter is close to the 1 and the second vanishes. Because the proof of Theorem 1 extends to any arbitrary fixed $c^{*}$, which need not vary with $\delta$ [^17] Theorem 1 is written with flexible commitment postures to compare equilibrium to the commitment payoffs which vary with the discount rate., methods there extend naturally to fixed postures in $c^{*}(1,0)$.
Throughout the statement of Theorem 2, suppose $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ for some $c^{*}\in c^{*}(1,0)$. I make the additional nontriviality [^18] This assumption can be derived by putting mild regularity conditions on $\kappa$, for example that $\kappa^{\prime}(0)=0$. assumption that $c^{*}\neq 0$, so that the commitment type is distinguishable from one playing static Nash.
@@
For all $\varepsilon>0$, there exists $\bar{\tau}>0$, $\bar{\delta}<1$, and $\bar{\varsigma}(\tau,\delta)$ vanishing as $\tau\to 0,\delta\to 1$, such that if $\text{var}(P_{\theta})=\tau<\bar{\tau}$, $\delta>\bar{\delta}$, and $\varsigma<\bar{\varsigma}(\tau,\delta)$, then
@@
    If $R(\mu)<0$, asymptotic $c^{*}$ incentive compatibility holds, $\underline{U}_{0}^{M}(\delta)\geq V^{*}(\delta)-\varepsilon$ and $f(a^{*}(c^{*}))>1-\varepsilon$.
@@
    If instead $R(\mu)>0$, the $\max\{\underline{U}_{0}(\delta),\bar{U}^{M}(\delta)\}<\varepsilon$.
Theorem 2 highlights a stark difference between the regime’s Stackelberg payoff depending on where their average regime strength when (1) the average becomes an arbitrarily precise summary of the state in each period, (2) the regime becomes arbitrarily patient, and finally (3) the global games perturbation vanishes. Importantly, it delivers a novel implication of risk dominance that is absent from the static game: the regime can only successfully deter rebellion if and only if they are, on average, able to deter rebellion at the risk-dominant level. Should they not be able to do this, then exit immediately happens with high probability.
The result stands in stark contrast to two seminal papers in the literature on dynamic global games with exit. First, it augments the model of Angeletos et al. (2007) to (1) small i.i.d. shocks to the state $\mu$ and (2) reputational incentives for the regime and shows how these perturbations can select a version of the risk-dominant equilibrium (and survival only if $\mu>\frac{1}{1+c^{*}}$) as the unique Markovian equilibrium. Finally it augments the reputational global game model of Huang (2017) to allow for small i.i.d. shocks and a rich exit process, and shows how the regime can fail if their average strength is too weak. This stands in stark contrast to his main result, where there are no attacks that ever occur on the equilibrium path. The differences in these results suggest that taking exit incentives seriously in reputational models can lead to rich results which are more indicative of real-world phenomena, where successful attacks do happen if speculator’s sense the regime is “likely weak.”
@@ 6 Discussion
This paper studied a model of reputation where players’ actions endogenously affected the probability of exit in the game (and hence the effective discount factor was endogenous). Such endogeneity creates several difficulties for the analysis of the repeated game. First, standard ways of normalizing payoffs no longer apply, and thus players’ values from different actions can vary directly with the discount rate. Second, folk theorems need no longer hold because the future can be “small” relative to actions today, making enforceability infeasible. Third, the addition of reputation effects need not select the long-run player’s Stackelberg payoff even when the long-run player becomes arbitrarily patient
I identify two conditions that help ameliorate these problems. These conditions are an incentive compatibility condition on the effective rate of future discounting: the long-run player wishes to play the Stackelberg action if recommended to do so, supposing maximal variation in continuation payoffs after; and a credible belief condition, represented by the Markovian refinement: short-run players believe the long-run player has an incentive to build a reputation when those incentives are present. Applied to a global games setting, these conditions yield new predictions about the survival of risk dominant equilibria in a dynamic regime change game.
There are several fruitful directions for future analysis. First, understanding the limits of endogenous exit when players’ actions are imperfectly monitored is a natural next step. The methods developed in this paper need not apply because reputation formation and incentive provision are much more subtle with imperfect monitoring, and the standard law-of-large-numbers intuitions do not apply with endogenous exit. Second, understanding the case with multiple commitment types is also promising, away from sequential equilibria—here, the set of payoffs may be pathological because of continuation payoffs at off-path deviations (i.e. taking a committed action that is not on the path of play). Understanding the set of Nash equilibria in this game and their payoff bounds is potentially promising. Finally, there is scope to apply the machinery in this paper to several other settings. For example, in the infinite chain store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982), my findings can give predictions about when the long-run player can deter entry even when entry affects the probability the incumbent firm must declare bankruptcy. Alternatively, Theorem 1 can be used to give conditions under which Markovian equilibria have reputation foundations in stopping time games where short-run or myopic players can stop at any time (i.e. adapted versions of Chassang (2010)). Finally, outside of Markovian predictions, these results may give new interpretations of “bad reputation” style results in the spirit of Ely and Valimaki (2003) and Ely et al. (2008).
@@
Proof of Lemma 1
Proof.
Let $\{\bar{f}_{\tilde{t}}\}\subset\{f_{t}\}$ be a subsequence such that $\lim\limits_{\tilde{t}\to\infty}\bar{f}_{\tilde{t}}=\bar{f}<1-\eta$, for some $\eta>0$ fixed; moreover, by truncating finitely many terms at most, suppose $\bar{f}_{\tilde{t}}<1-\eta$ for all $s\in\mathbb{N}$. Now for any $T\in\mathbb{N}$, define the number of times that $f_{t}<1-\eta$ to be $s(T)=\#\{s:f_{s}<1-\eta\}$. By assumption of $\eta$, we know $\lim\limits_{T\to\infty}s(T)=\infty$. We now have that $\prod_{s=0}^{\infty}f_{t}<(1-\eta)^{s(T)}$. Letting $\bar{v}$ be chosen so that $|v_{t}|\leq\bar{v}$ for all $t$, we have
$0\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}\delta^{t}v_{t}\prod_{s=0}^{\infty}f_{s}\right|\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}\delta^{t}\bar{v}(1-\eta)^{s(t)}\right|$
$\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{T}\delta^{t}\bar{v}(1-\eta)^{s(t)}\right|+\left|(1-\delta)(1-\eta)^{s(T)}\sum_{t=T+1}^{\infty}\delta^{t}\bar{v}\right|$
$\leq(1-\eta)^{s(T)}\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=T+1}^{\infty}\delta^{t}\bar{v}\leq(1-\eta)^{s(T)}\bar{v}.$
Proof of Lemma 2
Proof.
Note
$\sum_{t=0}^{\infty}\delta_{1}^{t}b_{t}=\delta^{T}\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}\geq 0\iff\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}$
@@
Some preliminaries. First, the limit exists since as $\delta$ increases the payoff set weakly expands. Moreover, $\underline{V}_{0}(\delta)$ must be the long-run player’s lowest possible equilibrium payoff as 0 is dominant in the stage game. Similarly, $V^{*}(\delta)$ must be their highest equilibrium payoff because short-run players must myopically best reply. Next, since $\mu_{0}(h^{t})=1$, we consider assessments where $\mu(h^{t})(\omega_{R})=1$ always and find subgame perfect equilibrium. Finally, throughout the proof, assume players have access to a public randomization device so that the payoff set is convex.
It is clear that $\underline{V}_{0}(\delta)$ is attainable by an equilibrium where the lowest stage-game Nash equilibrium is played at every history. Suppose now the second condition is satisfied. Consider the equilibrium with the following strategies.
@@
    At any other history, $\sigma_{L}^{*}(h^{t})=0$ and $\{\sigma_{i}(h^{t})\}_{i\in[N]}=\underline{a}\in\mathcal{N}(0)$, with $\underline{a}$ chosen to attain the lowest Nash equilibrium payoff $\underline{V}_{0}(\delta)$.
@@
$f(a^{*},c^{*})[v(a^{*},c^{*})+\delta V^{*}(\delta)]\geq f(a^{*},0)[v(a^{*},0)+\delta\underline{V}_{0}(\delta)].$
Divide through by $f(a^{*},0)$ and collect terms in $\delta$ on the left hand side to conclude that the above expression is equivalent to
$\delta[L(a^{*},c^{*})V^{*}(\delta)-\underline{V}_{0}(\delta)]\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*})$
as desired. Thus, so long as $\delta>\bar{\delta}$, $\underline{U}(\delta),\bar{U}(\delta)=[\underline{V}_{0}(\delta),V^{*}(\delta)]$.
@@
$f(a^{*},c^{*})[v(a^{*},c^{*})+\delta V(c^{*})]<f(a^{*},0)[v(a^{*},0)+\delta V(0)]$
and so it is never possible to induce the long-run player to play $c^{*}$ on path. Thus the long-run player’s largest payoff is bounded away from $\underline{V}_{0}^{*}(\delta)$ for any $\delta$. ∎
@@
Suppose not. Let $(\sigma,\mu)$ be an MPE and let $h^{t}$ be a history at which the long-run player mixes between 0 and $c^{*}$. By the argument in Lemmas 3 and 4, this can only possibly occur on a history $h^{t}=\{c^{*}\}_{s=0}^{t-1}$ for some $t$; call such a history the time-$t$ clean history. Define
@@
First, $|\mathcal{T}|$ cannot be infinite. Suppose not, so there is a sequence of times $\{t_{n}\}_{n=1}^{\infty}\in\mathcal{T}$. Note that when the long-run player mixes, both (1) $h^{t_{n}}\in\mathbb{P}^{\sigma}$ where $h^{t_{n}}$ is clean, and (2) the long-run player’s action is statistically identifiable compared to the committed type. This implies by a concentration of measure argument (see Mailath and Samuelson (2006) Chapter 15) that $\lim\limits_{n\to\infty}\mu(h^{t_{n}})(\omega_{c^{*}})=1$. Because short-run player actions are upper hemi-continuous in $\mu(h^{t_{n}})$, there must exist $N$ sufficiently large that for all $n\geq N$, $\mathcal{V}(h^{t_{n}}|\sigma)\geq\mathcal{V}(c^{*},h^{t_{n}})\geq V_{*}(\delta)-\varepsilon$ for some $\varepsilon>0$, where $\mathcal{V}(c^{*},h^{t})$ is the long-run player’s continuation history from simply always taking $c$ in the future and $\mathcal{V}(h^{t}|\sigma)$ is their continuation value from $\sigma$. Yet mixing implies the long-run player must be indifferent between taking 0 (and obtaining at most $\bar{V}_{0}(\delta)$ hereafter) and taking $c^{*}$ (and obtaining that future continuation value). In particular, take $N$ sufficiently large and $a=\{\sigma_{i}(h^{t_{N}})\}$ that
$f(a,0)[v(a,0)+\delta\mathcal{V}(\{h^{t},0\}|\sigma)]=f(a,c^{*})[v(a,c^{*})+\delta V(\{h^{t_{n}},c^{*}\}|\sigma)].$
@@
$f(a,0)[v(a,0)+\delta\mathcal{V}(\{h^{t},0\}|\sigma)]\leq f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]$
@@
$f(a,c^{*})[v(a,c^{*})+\delta V(\{h^{t_{n}},c^{*}\}|\sigma)]\leq f(a,c^{*})[v(a,c^{*})+\delta(V_{*}(\delta)-\varepsilon)].$
Yet the assumption the long-run player asymptotically satisfies $c^{*}$ IC implies
$f(a,0)[v(a,0)+\delta\bar{V}(\delta)]<f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
for $a$ since $a\in\mathcal{N}(\alpha c)$. Taking everything together and using asymptotic $c^{*}$-incentive compatibility implies for all $N$ sufficiently large that
$f(a,0)[v(a,0)+\delta\mathcal{V}(\{h^{t},0\}|\sigma)]<f(a,c^{*})[v(a,c^{*})+\delta V(\{h^{t_{n}},c^{*}\}|\sigma)]$
@@
Thus, $|\mathcal{T}|$ is finite. Let $T>0$ be the last time at which the long-run player mixes at a clean history. This implies after $T$, the long-run player attains at least $V_{*}(\delta)$ when they play 0. Thus, for $a=\{\sigma_{i}(h^{T})\}\in\mathcal{N}(\alpha c)$ for some $\alpha\in[0,1]$, we have
$f(a,0)[v(a,0)+\delta\underline{V}_{0}(\delta)]=f(a,c^{*})[v(a,c^{*})+\delta V^{*}(\delta)]$
This again contradicts asymptotic $c^{*}$ IC. Hence the long-run player never mixes. ∎
@@
Suppose the condition fails and fix any $\delta\in(0,1)$. There are three cases. First, suppose the inequality fails strictly at every $\alpha\in[0,1]$. Then in particular
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]>f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
This incentive compatibility constraint implies that the rational long-run player would always prefer to separate and play 0 (starting from any time-0 belief) if they are guaranteed $\bar{V}_{0}(\delta)$ afterwards. In particular, it implies the following is a Markovian equilibrium:
@@
    short-run players play $a_{*}=a^{*}(c^{*})$ at all $h^{t}$ where $\mu(h^{t})(\omega_{R})=0$, with $a_{*}$ chosen to deliver the long-run player expected continuation payoff $V_{*}(\delta)$.
@@
    If $\mu(h^{t})(\omega_{R})=1$, play $a_{0}^{*}\in\mathcal{N}(0)$ chosen to deliver the long-run player an expected continuation payoff $\bar{V}_{0}(\delta)$.
@@
Second, suppose instead that the inequality fails for some $\alpha$ but that it holds for $\alpha=1$. This implies first that it is possible to sustain the Markovian equilibrium where the long-run player plays $c^{*}$ at all histories. Moreover, there must exist some $\bar{\alpha}\in(0,1)$ with $a\in\mathcal{N}(\alpha c^{*})$ such that
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]=f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
by the intermediate value theorem, since the incentive compatibility constraint is continuous in $\alpha$. This implies that there exists a Markovian equilibrium of the following form:
@@
    If $\mu(h^{t})(\omega_{R})\neq\{\mu_{0}(\omega_{R}),1\}$, then have $\sigma_{L}(h^{t})=c^{*}$ and $\{\sigma_{i}(h^{t})\}\in\mathcal{N}(c^{*})$ be chosen to give the long-run player $V_{*}(\delta)$.
@@
    If $\mu(h^{t})(\omega_{R})$, play static Nash $\sigma(h^{t})$ giving the long-run player $\bar{V}_{0}(\delta)$.
@@
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]<f(a_{*},0)[v(a_{*},0)[v(a_{*},0)+\delta\bar{V}_{0}(\delta)]]$
$\leq f(a_{*},c^{*})[v(a_{*},c^{*})+\delta V_{*}(\delta)]$
where the first inequality follows from monotonicity of the game and the second follows from the fact the incentive compatibility condition holds at $\alpha=1$ where $a_{*}=a^{*}(c^{*})$ is the largest element of $N(c^{*})$, since by the monotonicity assumption this is the one which minimizes the stage-game payoff. Taking $\delta\to 1$ implies that the lowest equilibrium payoff is still separated from the commitment payoff by at least $f(a_{*},0)v(a_{*},0)-f(a,0)v(a,0)>0$; since this gap is independent of $\delta$, it implies the result.
Finally, suppose the inequality fails when $\alpha=1$. Then there cannot be an equilibrium where the long-run player imitates $c^{*}$ for sure. The only other pure Markovian equilibrium gives the long-run player their minimax payoff and implies the result. If instead the long-run player mixes in a Markovian equilibrium at some on-path history, they secure (for some $a\in\mathcal{N}(\alpha c^{*})$ and some $\alpha\in(0,1)$) a payoff of at most
$f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]<f(a^{*},c^{*})[v(a^{*},c^{*})+\delta V_{*}(\delta)]=V_{*}(\delta)$
by monotonicity and hence must secure a payoff smaller than $V_{*}(\delta)$. Taking $\delta\to 1$ finishes this last case and thus the proof. ∎

### v5

@@ 5.2 Selecting Equilibria
When is it that the threat of punishment is enough to deter revolution? Since the game is monotone and satisfies the conditions of Corollary 1, $c^{*}$-incentive compatibility is both necessary and sufficient for the regime to deter revolution (as much as possible) in the unique Markov equilibrium. Rewriting the incentive constraints to the simpler setting of the global game described above yields that $c^{*}$-incentive compatibility is equivalent to the inequality
$V_{*}-\underline{V}_{0}>\kappa(c^{*}).$
In the global games setting, $c^{*}$ incentive compatibility is related to the degree of dispersion of higher order beliefs around the risk-dominant threshold at the average, i.e. whether $\mu>(<)\frac{1}{1+c^{*}}$. Here, recall value $\frac{1}{1+c^{*}}$ is the risk-dominant signal threshold selected by the global game as signals become arbitrarily precise, i.e. $\lim\limits_{\varsigma\to 0}x^{*}(c^{*})=\frac{1}{1+c^{*}}$ by Frankel et al. (2003).
What is the right commitment posture $c^{*}$ to consider when comparing to $\mu$? There are two concerns. First, since $c^{*}(\varsigma)$ varies with the noise precision, the risk-dominant selection threshold varies with the noise precision as well. Second, the underlying probability distributions $(F_{\varepsilon},P_{\theta})$ will also affect the optimal choice for the regime and affect their commitment payoff. To resolve this problem, consider the set
$c^{*}(0)=\limsup_{\varsigma\to 0}c^{*}(\delta,\varsigma)$
where $\limsup A_{n}$ are the subsequential limits of all selections from $A_{n}$. That $c^{*}(0)$ is nonempty follows from a compactness arguments, noting $c^{*}(\varsigma)$ is well-defined and nonempty for each $\varsigma$ and hence so must $c^{*}(0)$. Throughout, I will fix some $c^{*}\in c^{*}(0)$ and appeal to continuity of the problem in parameters to make the argument that the same statement holds for $c^{*}(\varsigma)$.
Throughout the statement of Theorem 2, suppose $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ for some $c^{*}\in c^{*}(0)$. I also assume by fiat that $c^{*}\neq 0$, which is assumed throughout the main exposition, so that the commitment type is distinguishable from one playing static Nash. [^17] This assumption can be derived by putting mild regularity conditions on $\kappa$, for example that $\kappa^{\prime}(0)=0$.
@@
For all $\varepsilon>0$, there exists $\overline{\tau}>0$ and $\overline{\varsigma}(\tau)$ vanishing as $\tau\to 0$, such that if $\text{var}(\theta)=\tau<\overline{\tau}$, and $\varsigma<\overline{\varsigma}(\tau)$, then
@@
    If $\frac{1}{1+c^{*}}<\mu$, $\underline{U}_{0}^{M}=V^{*}$ and $f(a^{*}(c^{*}))>1-\varepsilon$.
@@
    If instead $\frac{1}{1+c^{*}}>\mu$, the $\max\{\underline{U}_{0},\overline{U}^{M}\}<\varepsilon$.
Theorem 2 highlights a stark difference between the regime’s Stackelberg payoff depending on where their average regime strength when (1) the average becomes an arbitrarily precise summary of the state in each period, (2) the regime becomes extremely patient, and finally (3) the global games perturbation vanishes. Importantly, it delivers a novel implication of risk dominance that is absent from the static game: the regime can only successfully deter rebellion if and only if they are, on average, able to deter rebellion at the risk-dominant level. Should they not be able to do this, then exit immediately happens with high probability.
The result stands in contrast to two seminal papers in the literature on dynamic global games with exit. First, it augments the model of Angeletos et al. (2007) to (1) small i.i.d. shocks to the state $\mu$ and (2) reputational incentives for the regime and shows how these perturbations can select a version of the “survival dominant” equilibrium (i.e. survival only if $\mu>\frac{1}{1+c^{*}}$) as the unique Markov equilibrium. Finally it augments the reputational global game model of Huang (2017) to allow for small i.i.d. shocks and a rich exit process, and shows how the regime can fail if their average strength is too weak. This stands in stark contrast to his main result, where there are no attacks that ever occur on the equilibrium path. The differences in these results suggest that taking exit incentives seriously in reputational models can lead to rich results which are more indicative of real-world phenomena, where successful attacks do happen if speculator’s sense the regime is “likely weak.”
@@ 6 Discussion
This paper studied a model of reputation where players’ actions endogenously affected the probability of exit in the game (and hence the effective discount factor was endogenous). Such endogeneity creates several difficulties for the analysis of the repeated game. First, folk theorems need no longer hold because the future can be “small” relative to actions today, making enforceability infeasible. Second, the addition of reputation effects need not select the long-run player’s Stackelberg payoff even when the long-run player becomes extremely patient.
I identify two conditions that help ameliorate these problems. These conditions are an incentive compatibility condition on the effective rate of future discounting: the long-run player wishes to play the Stackelberg action if recommended to do so, supposing maximal variation in continuation payoffs after; and a credible belief condition, represented by the Markov refinement: short-run players believe the long-run player has an incentive to build a reputation when those incentives are present. Applied to a global games setting, these conditions yield new predictions about the survival of equilibria relative to the risk dominant threshold in a dynamic regime change game.
There are several fruitful directions for future analysis. First, understanding the limits of endogenous exit when players’ actions are imperfectly monitored is a natural next step. The methods developed in this paper need not apply because reputation formation and incentive provision are much more subtle with imperfect monitoring, and the standard law-of-large-numbers intuitions do not apply with endogenous exit. Second, understanding the case with multiple commitment types is also promising, away from sequential equilibria—here, the set of payoffs may be pathological because of continuation payoffs at off-path deviations (i.e. taking a committed action that is not on the path of play). Understanding the set of Nash equilibria in this game and their payoff bounds is potentially promising. Finally, there is scope to apply the machinery in this paper to several other settings. For example, in the repeated trust game, my findings can give predictions about when the long-run player can teach short-run players to trust them even when a lack of trust might prematurely end the game. Alternatively, Theorem 1 can be used to give conditions under which Markov equilibria have reputation foundations in stopping time games where short-run or myopic players can stop at any time (i.e. adapted versions of Chassang (2010)). Finally, outside of Markov predictions, these results may give new interpretations of “bad reputation” style results in the spirit of Ely and Valimaki (2003) and Ely et al. (2008).
@@
Some preliminaries. First, $\underline{V}_{0}$ must be the long-run player’s lowest possible equilibrium payoff as 0 is dominant in the stage game. Similarly, $V^{*}$ must be their highest equilibrium payoff because short-run players must myopically best reply. Next, since $\mu_{0}(h^{t})=1$, we consider assessments where $\mu(h^{t})(\omega_{R})=1$ always and find subgame perfect equilibrium. Finally, throughout the proof, assume players have access to a public randomization device so that the payoff set is convex.
It is clear that $\underline{V}_{0}$ is attainable by an equilibrium where the lowest stage-game Nash equilibrium is played at every history. Suppose now the second condition is satisfied. Consider the equilibrium with the following strategies.
@@
    At any other history, $\sigma_{L}^{*}(h^{t})=0$ and $\{\sigma_{i}(h^{t})\}_{i\in[N]}=\underline{a}\in\mathcal{N}(0)$, with $\underline{a}$ chosen to attain the lowest Nash equilibrium payoff $\underline{V}_{0}$.
@@
$f(a^{*},c^{*})[v(a^{*},c^{*})+V^{*}]\geq f(a^{*},0)[v(a^{*},0)+\underline{V}_{0}].$
Divide through by $f(a^{*},0)$ and collect terms on the left hand side to conclude that the above expression is equivalent to
$(V^{*})L(a^{*},c^{*})-\underline{V}_{0}\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*})$
as desired. Thus, $\underline{U},\overline{U}]=[\underline{V}_{0},V^{*}]$.
@@
$f(a^{*},c^{*})[v(a^{*},c^{*})+V(c^{*})]<f(a^{*},0)[v(a^{*},0)+V(0)]$
and so it is never possible to induce the long-run player to play $c^{*}$ on path. Thus the long-run player’s largest payoff is bounded away from $\underline{V}_{*}$. Note this argument holds even for the discounted repeated game, for any discount rate. ∎
@@
Suppose not. Let $(\sigma,\mu)$ be an MPE and let $h^{t}$ be a history at which the long-run player mixes between 0 and $c^{*}$. By the argument in Lemmas 1 and 2, this can only possibly occur on a history $h^{t}=\{c^{*}\}_{s=0}^{t-1}$ for some $t$; call such a history the time-$t$ clean history. Define
@@
First, $|\mathcal{T}|$ cannot be infinite. Suppose not, so there is a sequence of times $\{t_{n}\}_{n=1}^{\infty}\in\mathcal{T}$. Note that when the long-run player mixes, both (1) $h^{t_{n}}\in\mathbb{P}^{\sigma}$ where $h^{t_{n}}$ is clean, and (2) the long-run player’s action is statistically identifiable compared to the committed type. This implies by a concentration of measure argument (see Mailath and Samuelson (2006) Chapter 15) that $\lim\limits_{n\to\infty}\mu(h^{t_{n}})(\omega_{c^{*}})=1$. Because short-run player actions are upper hemi-continuous in $\mu(h^{t_{n}})$, there must exist $N$ sufficiently large that for all $n\geq N$, $\mathcal{V}(h^{t_{n}}|\sigma)\geq\mathcal{V}(c^{*},h^{t_{n}})\geq V_{*}-\varepsilon$ for some $\varepsilon>0$, where $\mathcal{V}(c^{*},h^{t})$ is the long-run player’s continuation history from simply always taking $c$ in the future and $\mathcal{V}(h^{t}|\sigma)$ is their continuation value from $\sigma$. Yet mixing implies the long-run player must be indifferent between taking 0 (and obtaining at most $\overline{V}_{0}$ hereafter, since they have revealed their rationality) and taking $c^{*}$ (and obtaining some other future continuation value). In particular, for $N$ sufficiently large and $a=\{\sigma_{i}(h^{t_{N}})\}$, we have
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t},0\}|\sigma)]=f(a,c^{*})[v(a,c^{*})+V(\{h^{t_{n}},c^{*}\}|\sigma)].$
@@
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t},0\}|\sigma)]\leq f(a,0)[v(a,0)+\overline{V}_{0}]$
@@
$f(a,c^{*})[v(a,c^{*})+V(\{h^{t_{n}},c^{*}\}|\sigma)]\geq f(a,c^{*})[v(a,c^{*})+(V_{*}-\varepsilon)].$
Yet the assumption the long-run player satisfies $c^{*}$ IC implies
$f(a,0)[v(a,0)+\overline{V}_{0}]<f(a,c^{*})[v(a,c^{*})+V_{*}]$
for $a$ since $a\in\mathcal{N}(\alpha c)$. Taking everything together and using $c^{*}$-incentive compatibility implies for all $N$ sufficiently large that
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t},0\}|\sigma)]<f(a,c^{*})[v(a,c^{*})+V(\{h^{t_{n}},c^{*}\}|\sigma)]$
@@
Thus, $|\mathcal{T}|$ is finite. Let $T>0$ be the last time at which the long-run player mixes at a clean history. This implies after $T$, the long-run player attains at least $V_{*}$ when they play 0. Thus, for $a=\{\sigma_{i}(h^{T})\}\in\mathcal{N}(\alpha c)$ for some $\alpha\in[0,1]$, we have
$f(a,0)[v(a,0)+\underline{V}_{0}]=f(a,c^{*})[v(a,c^{*})+V^{*}]$
This again contradicts $c^{*}$ IC. Hence the long-run player never mixes. ∎
@@
Suppose the condition fails. There are three cases. First, suppose the inequality fails strictly at $\alpha=0$. Then in particular
$f(a,0)[v(a,0)+\overline{V}_{0}]>f(a,c^{*})[v(a,c^{*})+V_{*}]$
for $a\in\mathcal{N}(0)$. This incentive compatibility constraint implies that the rational long-run player would always prefer to separate and play 0 (starting from any time-0 belief) if they are guaranteed $\overline{V}_{0}$ afterwards. In particular, it implies the following is a Markov equilibrium as $\mu_{0}(\omega_{R})\to 1$:
@@
    Short-run players play $a_{*}=a^{*}(c^{*})$ at all $h^{t}$ where $\mu(h^{t})(\omega_{R})=0$, with $a_{*}$ chosen to deliver the long-run player expected continuation payoff $V_{*}$.
@@
    If $\mu(h^{t})(\omega_{R})=1$, play $a_{0}^{*}\in\mathcal{N}(0)$ chosen to deliver the long-run player an expected continuation payoff $\overline{V}_{0}$.
@@
Second, suppose instead that the inequality fails for some $\alpha\in(0,1)$ but that it holds for $\alpha=1$. This implies first that it is possible to sustain the Markov equilibrium where the long-run player plays $c^{*}$ at all histories. Moreover, there must exist some $\overline{\alpha}\in(0,1)$ with $a\in\mathcal{N}(\alpha c^{*})$ and
$f(a,0)[v(a,0)+\overline{V}_{0}]=f(a,c^{*})[v(a,c^{*})+V_{*}]$
by the intermediate value theorem, since the incentive compatibility constraint is continuous in $\alpha$. This implies that there exists a Markov equilibrium of the following form:
@@
    If $\mu(h^{t})(\omega_{R})\neq\{\mu_{0}(\omega_{R}),1\}$, then have $\sigma_{L}(h^{t})=c^{*}$ and $\{\sigma_{i}(h^{t})\}\in\mathcal{N}(c^{*})$ be chosen to give the long-run player $V_{*}$.
@@
    If $\mu(h^{t})(\omega_{R})$, play static Nash $\sigma(h^{t})$ giving the long-run player $\overline{V}_{0}$.
@@
$f(a,0)[v(a,0)+\overline{V}_{0}]<f(a_{*},0)[v(a_{*},0)[v(a_{*},0)+\overline{V}_{0}]]$
$\leq f(a_{*},c^{*})[v(a_{*},c^{*})+V_{*}]$
where the first inequality follows from monotonicity of the game and the second follows from the fact the incentive compatibility condition holds at $\alpha=1$ where $a_{*}=a^{*}(c^{*})$ is the largest element of $\mathcal{N}(c^{*})$, since again by the monotonicity assumption this is the choice which minimizes the stage-game payoff. This implies the result.
Finally, suppose the inequality fails when $\alpha=1$. Then there cannot be an equilibrium where the long-run player imitates $c^{*}$ for sure. The only other pure Markov equilibrium gives the long-run player their minimax payoff and implies the result. If instead the long-run player mixes in a Markov equilibrium at some on-path history, they secure (for some $a\in\mathcal{N}(\alpha c^{*})$ and some $\alpha\in(0,1)$) a payoff of at most
$f(a,c^{*})[v(a,c^{*})+V_{*}]<f(a^{*},c^{*})[v(a^{*},c^{*})+V_{*}]=V_{*}$
by monotonicity and hence must secure a payoff smaller than $V_{*}$. ∎